Expected return A stocks returns have the following distribution: Demand for the Companys Products Weak Below average Average Above average Strong Rate of Return If Probability of This Demand Occurring 0.2 0.1 0.4 0.2 0.1 1.0 This Demand Occurs -50% -10 17 20 49 a. Calculate the stocks expected return. Round your answer to two decimal places. b. Calculate the stocks standard deviation. Round your answer to two decimal places. c. Calculate the stocks coefficient of variation. Round your answer to two decimal places.if using excel please show formulas so i know how to do.
Answer #1

Expected return=Respective return*Respective probability

=(0.2*-50)+(0.1*-10)+(0.4*17)+(0.2*20)+(0.1*49)=4.7%

Probability REturn Probability*(Return-Mean)^2
0.2 -50 0.2*(-50-4.7)^2=598.418
0.1 -10 0.1(-10-4.7)^2=21.609
0.4 17 0.4*(17-4.7)^2=60.516
0.2 20 0.2*(20-4.7)^2=46.818
0.1 49 0.1*(49-4.7)^2=196.249
Total=923.61

Standard deviation=[Total Probability*(Return-Mean)^2/Total Probability][^(1/2)

=30.39%(Approx)

Coefficient of variation=Standard deviation/mean

=(30.39/4.7)=6.47(Approx).